Friendly’s Quick Loans, Inc., offers you “six for eight or I knock on your door.” This means you get $6 today and repay $8 when you get your paycheck in one week (or else).

what is the APR and EAR
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

8=6*(1+r/100)

(8/6)=(1+r/100)

(1+r/100)=1.3333

r=1.3333-1

=33.33%(Approx)

APR=33.33%/week*52 weeks

=1733.33%(Approx)

EAR=[(1+APR/m)^m]-1
where m=compounding periods

=[(1+0.3333)^52]-1

=313,916,515.7%(Approx).